OpenStudy (friendzone):

: -2/3 + c = 2/3 * Be sure to solve for c by balancing the equation!

OpenStudy (immanuelv):

you could move the -2/3 to the other side and you'll get c

OpenStudy (immanuelv):

c = 2/3 + 2/3

undeadknight26 (undeadknight26):

Although when we check out work its wrong ._. -2/3 + c = 2/3 -2/3 + 0 = 2/3? You smells dat ._. Something looks fishy ._. Or i could just be tired and seeing things ._.